8 coins are tossed simultaneously. The probability of getting at least 6 heads is
- AOption (a)
7/64
- BOption (b)
57/64
- COption (c)
37/256
- DOption (d)
229/256
View answer & explanationHide answer & explanation
Correct answer
37/256
Why this answer is correct
Concept:
- Probability of exactly r success in n trials \( = {\rm{;p}}\left( {{\rm{x}} = {\rm{r}}} \right) = {\rm{;}}{{\rm{;}}^{\rm{n}}}{{\rm{C}}_{\rm{r}}}{\rm{;}}{{\rm{p}}^{\rm{r}}}{{\rm{q}}^{{\rm{n}} - {\rm{r}}}}\)
Where n = number of trials; p = probability of success; q = (1-p) = probability of failure
Calculation:
We have to find the probability of getting at least 6 heads
Given, n = 8
Here Success is getting a head on single trial.
We already know that there are exactly 2possible outcomes for each and every toss of coins
Probability of success = p = 1/2
Probability of failure = q = 1 – p = 1 – 1/2 = 1/2
The probability of getting at least 6 heads = P (x = 6) + P (x = 7) + P (x = 8)
\( = {{\rm{;}}^8}{{\rm{C}}_6}{\rm{;}}{\left( {\frac{1}{2}} \right)^6}{\left( {\frac{1}{2}} \right)^{8 - 6}} + {{\rm{;}}^8}{{\rm{C}}_7}{\rm{;}}{\left( {\frac{1}{2}} \right)^7}{\left( {\frac{1}{2}} \right)^{8 - 7}} + {{\rm{;}}^8}{{\rm{C}}_8}{\rm{;}}{\left( {\frac{1}{2}} \right)^8}{\left( {\frac{1}{2}} \right)^{8 - 8}}\)
\( = ;{\left( {\frac{1}{2}} \right)^8};\left( {{{\rm{;}}^8}{{\rm{C}}_6}{\rm{;}} + {{\rm{;}}^8}{{\rm{C}}_7}{\rm{;}} + {{\rm{;}}^8}{{\rm{C}}_8}{\rm{;}}} \right) = ;{\left( {\frac{1}{2}} \right)^8}\left( {28 + 8 + 1} \right) = ;\frac{{37}}{{256}}\)